# What To Do (and Not to Do) with Causal Panel Analysis under Parallel Trends:

### v2

@@ 4.  Data and Procedure
Finally, we conduct diagnostic tests based on FEct to further assess the PTA and, in applications with treatment reversal, the no-carryover-effect assumption. We use FEct because it is applicable across all studies in our replication sample and does not discard data. One important caveat is that the results of these formal statistical tests need to be combined with visual inspections of DTE plots, as they may either suffer from low power with a small sample size or be “too powerful” with abundant data. [^11] By a test being too powerful, we mean that it has power against alternatives that are very close to the null (which might occur, for example, in the presence of an outlier or a marginally consequential confounder) and unlikely to be a source of significant bias in the treatment effect estimates. To guard against the latter possibility, we also perform an equivalence test of the null that pre-treatment (or carryover) effects are “large,” i.e. outside of some $\theta$-neighborhood of 0. Note this test requires us to specify a threshold $\theta$, and we conduct tests with a default threshold and with the ATT set as the threshold. [^12] We provide the implementation details of all tests mentioned in this section in the ‘Diagnostics Tests’ section of the SM and refer readers to Liu, Wang and Xu (2022) for more details on all the tests we use, including how $\theta$ is chosen. Small $p$-values are evidence in favor of the claim that PTA violations are either nonexistent or mild. Figures 3(d), (e) and (f) show the results from the $F$ test, placebo test, and test for no carryover effects on our running example, respectively. Both a visual inspection and the formal tests suggest that the PTA and the no-carryover-effect assumption are quite plausible.
@@
On the whole, the result from Grumbach and Sahn (2020) that we reanalyze seems highly robust. The study also seems to have sufficient power to not only distinguish the ATT from zero, but also to assess the validity of its identification assumptions. There is strong evidence that the PTA is satisfied and that there are no carryover effects. The results are consistent across choice of estimator and robust to using cluster-bootstrapped standard errors.
@@ 5.  Systematic Assessment
We carry out the replication and reanalysis procedure described above for all 37 articles in our sample. This section offers a summary of our findings; the complete results for each paper can be found in the SM. We structure our results around two main questions: (1) Is the PTA plausible, and (2) do HTE-robust estimators provide results that are qualitatively different from those obtained with TWFE? Additionally, we report on other issues we observe in the replicated articles, which include lack of statistical power, presence of carryover effects, and lack of robustness to alternative specifications, such as the addition of LDVs or ULTs.
PTA violations remain a concern.
Although the recent methodological literature heavily focuses on the issue of HTE, we report that PTA violations remain a major concern in practice despite having been a well-known pitfall for some time now. Notably, over half of the studies in our sample do not include a DTE/event-study plot of any kind in the main text or SM. In Figure 4, we plot the DTE estimates from FEct. We also report the ATT estimates and their bootstrapped SEs, as well as the $p$-values of the $F$ tests and equivalence tests used to assess the presence of “pretrends.” Due to space limitations, we present the DTE plots from other estimators, as well as results from the placebo tests, in the SM.
@@
$p$-values: 0.19, 0.71 Bischof and Wagner (2019)
@@
$p$-values: 0.79, 0.49 Bisgaard and Slothuus (2018)
@@
ATT: 0.10 (0.03);
$p$-values: 0.04, 0.90 Caughey, Warshaw and Xu (2017)
@@
$p$-values: 0.32, 0.51 Christensen and Garfias (2021)
@@
$p$-values: 0.77, 0.30 Clarke (2020)
ATT: 0.10 (0.02);
$p$-values: 0.29, 0.00
@@
$p$-values: 0.00, 0.00 Distelhorst and Locke (2018)
@@
$p$-values: 0.15, 0.00 Fouirnaies and Hall (2022)
@@
$p$-values: 0.59, 0.42 Grumbach and Sahn (2020)
ATT: 0.14 (0.03);
$p$-values: 0.07, 0.00 Grumbach and Hill (2022)
@@
$p$-values: 0.68, 0.69 Hainmueller and Hangartner (2019)
@@
$p$-values: 0.36, 0.00 Kilborn and Vishwanath (2022)
ATT: 0.05 (0.02);
@@
$p$-values: 0.86, 0.00 Magaloni, Franco-Vivanco and Melo (2020)
@@
$p$-values: 0.87, 0.02 Pierskalla and Sacks (2018)
@@
$p$-values: 0.65, 0.00 Ravanilla, Sexton and Haim (2022)
@@
Refer to caption Note: We report the estimated ATT and corresponding bootstrap SEs (in parentheses) using FEct. We also provide the $p$-values from the $F$ test and equivalence test. Rejecting the null in the $F$ test indicates potential violations of the PTA while rejecting the null in the equivalence test provides evidence in support of the PTA. These tests are infeasible for four cases with only one pre-treatment period.

### v3

@@ 4.  Data and Procedure
One important caveat is that the results of these tests may either suffer from low power with a small sample size or be “too powerful” with abundant data. [^14] By a test being too powerful, we mean that it has power against alternatives that are very close to the null (which might occur, for example, in the presence of an outlier or a marginally consequential confounder) and unlikely to be a source of significant bias in the treatment effect estimates. To address the latter concern, we also perform an equivalence test of the null that pretreatment effects are “large,” i.e. outside some $\theta$-neighborhood of 0. Note this test requires us to specify a threshold $\theta$, and we conduct tests with a default threshold and with the ATT set as the threshold. [^15] We provide the implementation details of all tests mentioned in this section in Section A.4. of the SM and refer readers to Liu, Wang and Xu (2022) for more details on all the tests we use, including how $\theta$ is chosen. Small $p$-values are evidence in favor of the claim that PTA violations are at most mild.
Finally, we compute robust CSs proposed by Rambachan and Roth (2023) that allow for PTA violations under relative magnitude (RM) restrictions, which allow for post-treatment violations of the PTA that are at most $\bar{M}$ times the size of the maximum violation in pretreatment placebo periods. [^16] We use pretreatment placebo periods as the benchmark because placebo estimates are estimated in the same way as post-treatment ATT estimates (Roth, 2024). In other words, the differential trend between any two consecutive post-treatment periods are assumed to be no greater than $\bar{M}$ times the maximum difference between any two placebo periods. If we first assume we can decompose each estimator for the DTE $\beta_{t}$ into the actual DTE $\tau_{t}$ and a trend component $\delta_{t}$, $\beta_{t}=\tau_{t}+\delta_{t}$, then the RM assumption requires that for all periods relative to treatment $t$,
$|\delta_{t+1}-\delta_{t}|\leq\bar{M}\cdot\max_{s\in\mathcal{P}\backslash\{0\}}\{\delta_{s+1}-\delta_{s}\}$.
in which $\mathcal{P}$ is the set of placebo periods. In our case, we use $\mathcal{P}=\{-2,-1,0\}$, meaning the maximum trend is given by $\max\{\delta_{0}-\delta_{-1},\delta_{-1}-\delta_{-2}\}$. Robust CSs are uniformly valid for the partially identified treatment effects under the RM restriction. Rambachan and Roth (2023, p. 2653) suggest $\bar{M}=1$ as a “natural benchmark” with similar numbers of pretreatment and post-treatment periods, corresponding to the assumption that any PTA violations are no worse than the observed pretrend. For each paper, we compute the robust CSs with $\bar{M}=0.5$, erring on the lenient side. We also conduct a sensitivity analysis for each such study, obtaining the “breakdown value” of $\bar{M}$, denoted as $\tilde{M}$, at which the robust CS first includes zero. Figure 2(e) marks the robust CS for the ATT estimated by FEct in pink. In this example, the robust CS excludes zero, and the breakdown value from the sensitivity analysis is $\tilde{M}=3.0$, suggesting that the estimated coethnic mobilization effect is robust to potential violations of the PTA under the RM restriction.
@@
Overall, the results from Grumbach and Sahn (2020) appear highly robust, regardless of the choice of point and variance estimators. The study seems to have sufficient power to evaluate and instill confidence in the validity of its identification assumptions and to distinguish the ATT from zero, even under violations of the PTA. There are also little to no carryover effects.
@@ 5.  Systematic Assessment
We perform the replication and reanalysis procedure described above for all 37 articles in our sample. This section offers a summary of our findings; complete results for each paper are available in the SM. We organize our results around two main questions: (1) Do HTE-robust estimators yield qualitatively different results compared to those obtained with TWFE? (2) Is the PTA plausible, and do original findings (where null hypotheses are rejected) remain robust against HTE and potential PTA violations? Additionally, we discuss other issues observed in the replicated articles, including the presence of carryover effects and sensitivity to model specifications.
HTE-robust estimators yield qualitatively similar but highly variable findings.
To examine the possible impact of the weighting problem caused by HTE, we first compare the estimates obtained from the imputation estimator, FEct, for all studies to those originally reported. The dissimilarity between the original and FEct estimates proxies the extent to which the negative weighting problem is consequential in practice. Figure 3(a) plots the comparison. The horizontal axis represents the originally reported TWFE estimates, and the vertical axis represents estimates from FEct, both normalized using the same originally reported SE. If the point estimates are identical, then the corresponding point should lie exactly on the 45-degree line (dashed). Red solid circles represent studies whose FEct estimates are not significant at the 5% level based on cluster-bootstrapped CIs. Observe that the points largely follow the 45-degree line, and that FEct always has the same sign as the original estimate. This suggests that scenarios where accounting for HTE entirely reverses the empirical findings, while theoretically possible, are rare.
This said, results do sometimes deviate significantly. In the extremes, we observe FEct estimates as small as a fourth of or as large as over twice the TWFE estimate. In the SM, we include a plot of the ratios of the estimates. The most consequential deviations correspond to studies that were originally on the margins of statistical significance. Alarmingly, when switching from TWFE to FEct, the number of studies that are statistically insignificant at the 5% level rises from one to 11.171717This is due to a combination of drops in point estimates and larger SEs. See Section A.5. in the SM for more details. This group constitutes 28% of all articles and disproportionately consists of studies whose original $z$ scores are smaller than three. We further discuss the consequences on power and inference later.
Figure 3: Comparison of Estimates: TWFE vs. HTE-Robust Estimators
(a) Comparison of FEct and TWFE estimates
(b) Comparison of HTE-robust and TWFE estimates in the staggered setting
Note: Subfigure (a) compares TWFE coefficients and estimates from the imputation method (FEct). Estimates for each application are normalized by the same TWFE SE. Fouirnaies and Hall (2018) and Hall and Yoder (2022) are close to the 45-degree line but not included in the figure as their TWFE $z$-scores are above 15. Subfigure (b) compares reported coefficients and estimates from various alternative estimators. All seven estimates for each application are normalized by the same reported SE. We multiply the estimates by $Sign(\hat{\delta}^{TWFE}$) for easier visualization. In both figures, red symbols represent studies whose ATT estimates from the respective HTE-robust estimators are statistically insignificant at the 5% level based on bootstrapped CIs.
If we restrict our attention to the staggered case, we can broaden our comparison set to include all HTE-robust estimators discussed earlier. [^18] Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are not included because the original specifications include additional time trends, which are not supported by HTE-robust estimators except FEct. Figure 3(b) plots the point estimates for these cases obtained from each HTE-robust estimator, in addition to TWFE, all normalized by the reported SE. The estimates from all HTE-robust estimators are qualitatively similar to each other, though there is a noticeable amount of variation and one sign change. As before, this spread does not change substantive findings for studies with highly substantively significant results. Furthermore, our more extensive analyses of each individual paper in the SM show that the DTE estimates from various HTE-robust estimators also generally align with each other and with those from the dynamic specification of the TWFE model.
It is worth noting that while only a small fraction of studies in our sample (five articles, 13.5%) employ a bootstrap procedure to estimate SEs or CIs, the widely practiced cluster-robust SE typically performs adequately because the number of clusters/units is usually fairly large. There are two examples that were significant at the $5\%$ level using cluster-robust SEs but fell below this threshold when using cluster-bootstrapped SEs, both of which were already on the margins of significance. We provide graphs comparing reported, cluster-robust, and cluster-bootstrapped SEs in the SM.
@@
$p$-values: 0.22, 0.68 Bischof and Wagner (2019)
@@
$p$-values: 0.79, 0.26 Bisgaard and Slothuus (2018)
@@
ATT: 0.11 (0.03);
$p$-values: 0.00, 0.86 Caughey, Warshaw and Xu (2017)
@@
$p$-values: 0.04, 0.52 Christensen and Garfias (2021)
@@
$p$-values: 0.72, 0.28 Clarke (2020)
ATT: 0.08 (0.03);
$p$-values: 0.38, 0.00
@@
$p$-values: 0.10, 0.00 Distelhorst and Locke (2018)
@@
$p$-values: 0.21, 0.00 Fouirnaies and Hall (2022)
@@
$p$-values: 0.90, 0.42 Grumbach and Sahn (2020)
ATT: 0.13 (0.03);
$p$-values: 0.42, 0.00 Grumbach and Hill (2022)
@@
$p$-values: 0.86, 0.37 Hainmueller and Hangartner (2019)
@@
$p$-values: 0.51, 0.00 Kilborn and Vishwanath (2022)
ATT: 0.12 (0.04);
@@
$p$-values: 0.85, 0.00 Magaloni, Franco-Vivanco and Melo (2020)
@@
$p$-values: 0.92, 0.03 Pierskalla and Sacks (2018)
@@
$p$-values: 0.01, 0.00 Ravanilla, Sexton and Haim (2022)
@@
Refer to caption Note: We report the estimated ATT and corresponding bootstrap SEs (in parentheses) using FEct. We also provide the $p$-values from the $F$ test and equivalence test. Rejecting the null in the $F$ test indicates potential violations of the PTA while rejecting the null in the equivalence test provides evidence in support of the PTA. These tests are infeasible for four cases with only one pretreatment period.
Allowing mild PTA violations causes most results to lose power.
Although the recent methodological literature heavily focuses on HTE, PTA violations remain a primary concern in practice, despite having been a well-known pitfall for some time now. Notably, over half of the studies in our sample do not include an event-study plot of any kind in the main text or appendices. In Figure 4, we present the event-study plot based on estimates from FEct for all papers in our sample. We also report the ATT estimates and their bootstrapped SEs, as well as the $p$-values of the $F$ tests and equivalence tests used to assess pretrends. Due to space limitations, we present the event-study plots from other estimators, as well as results from the placebo tests, robust CSs, and sensitivity analyses, in the SM.